See Conceptual Example 6 to review the concepts involved in this problem. A 11.2-kg monkey is hanging by one arm from a branch and swinging on a vertical circle. As an approximation, assume a radial distance of 67.6 cm is between the branch and the point where the monkey's mass is located. As the monkey swings through the lowest point on the circle, it has a speed of 2.11 m/s. Find (a) the magnitude of the centripetal force acting on the monkey and (b) the magnitude of the tension in the monkey's arm.
Answer #1

M = 11-2 lg R2 67-6cm = 0.676m V=2.11 m/s a) centripetal forre P fca 11:2x2-102 0.676 [Fe = 73.8N Part-B) 12 S Tension T = my
Answer #2

@ Given, m=11:2kg. vz 2.11mis. 8= 67.6CM = 0.676m centripetal farce is, Fez mu?= 112%2:11 0.676 ED) = 7376 N. Tea simo in hе